If you are expecting to receive an amount of $ 10,000 but after 2 years, how...

  1. If you are expecting to receive an amount of $ 10,000 but after 2 years, how much the present value would be today if the discount rate is 10%?
Answer #1

FV= Future value = 10000, YEARS = 2, RATE = 10%

we have to find PV = Present value

PV = FV/(1+r)^n = 10000/(1+0.10)^2 = 8264.46

Answer : 8264.46 (Thumbs up please)